Solve by factoring and using principle of zero products:
n^2+11n+18=0

Found 2 solutions by benn143, ewatrrr:
Answer by benn143(2)

n^2+11n+18=0
(n+9)(n+2)=0
n+9=0 n+2=0
n=-9 and n=-2

Answer by ewatrrr(24785)

 
Hi
Solve by factoring and using principle of zero products:
n^2+11n+18=0
(x+2)(x+9)= 0 Note:SUM of the outside terms(9n) and the inside terms(2n) = 11n
(x+2)= 0 x = -2
(x+9)= 0 x = -9
x = -9,-2